OpenStudy (anonymous):

Find the period, to the nearest tenth of a second, of a pendulum of length 4.8 feet.

OpenStudy (anonymous):

Do you know the formula for simple harmonic oscillation of a pendulum's period?

OpenStudy (anonymous):

no, i dont.

OpenStudy (anonymous):

Well, it's 2*pi*sqrt(L/g) where L = length of pendulum and g = gravity

OpenStudy (anonymous):

Ok, so how do I solve this problem?

OpenStudy (anonymous):

Just plug in the numbers...which is only the length of the pendulum

OpenStudy (anonymous):

is this correct 2*pi*sqrt(4.8) ?

OpenStudy (anonymous):

you have to divide the length by the force of gravity first (-9.8)

OpenStudy (anonymous):

i don't really get it. I'm not really familiar with the formula

OpenStudy (shubhamsrg):

and 4.8 wont be it..you'll have to convert it to meters ..since you're usong g=9.8 m/s

OpenStudy (anonymous):

Ok. \[T = 2\pi \sqrt{\frac{4.8}{9.8}}\]

OpenStudy (anonymous):

oh it says feet

OpenStudy (shubhamsrg):

1 foot= 0.305 meters

OpenStudy (anonymous):

we have to use length in feet and time in seconds

OpenStudy (anonymous):

is the answer 4.4?